OpenStudy (anonymous):

Integrate: logx/(1+logx)^2

OpenStudy (turingtest):

@mathguy1125 are those natural logs or log base ten?

hartnn (hartnn):

put u= 1+ log x so log x of numerator will be u-1. can u then solve it ?

OpenStudy (anonymous):

@Natural logs. @hartnn It has a slicker way, apparently. A substitution.

hartnn (hartnn):

hmmm.... i couldn't think of a substitution..... maybe e^x ??

OpenStudy (anonymous):

I dunno. Help? @TuringTest @sauravshakya @hartnn @experimentX

hartnn (hartnn):

didn't u try what i said ?

OpenStudy (anonymous):

You are saying: = 1/(1+logx) - logx/(1+logx)^2 Then?

hartnn (hartnn):

nopes, i was saying to put u= 1+log x dx = e^(u-1) du then integrate by parts..